OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the ellipse at the given point. (2,4) (x^2/4)+(y^2/8)=3 i got the equation to use -2x/y and then the equation after plugging in i got (y-4)=-1(x-2) not sure why thats wrong

OpenStudy (amistre64):

2/4 x + 2/8 y y' = 0 x/2 + y y'/4 = 0 y y'/4 = -x/2 y y' = -2x y' = -2x/y is good

OpenStudy (amistre64):

you didnt provide a point but lets use (a,b) the slope of the tangent line is -2a/b so we can use the point slope formula

OpenStudy (amistre64):

y = -2a(x-a)/b + b

OpenStudy (anonymous):

oh sorry the point is (2,4)

OpenStudy (anonymous):

i plug into the formula you gave me but says still wrong

OpenStudy (amistre64):

then there is something about the format that it is not liking. can you post a picture of the problem to verify you have posted the correct information?

OpenStudy (anonymous):

cant i just tell you what answer i got dont really have a camera on me

OpenStudy (amistre64):

is it on your screen? or on paper?

OpenStudy (anonymous):

and the info i posted is correct checked it twice

OpenStudy (amistre64):

when i work with a program like mathlab, i tend to read over some instructions that ask for a sepcific formating which is why i was wondering

OpenStudy (anonymous):

nope theres no asking for any type of formatting this is webassign they dont write on there in what form they want my professor told us to try the form i wrote out or slope intercept

OpenStudy (amistre64):

y = -(x-2) + 4 y = -x+6

OpenStudy (dumbcow):

you may want to consult webassign to find out if there any IT issues or something. if its consistently not allowing you to submit correct answers, thats not good

OpenStudy (anonymous):

alot of students complain about i but some reason even though its Purdue they wont take time to fix it have no idea why

OpenStudy (anonymous):

they give us 3 tries though because of the reason it takes correct answers it just has to be the form they wont they just dont ever post what form they want

OpenStudy (amistre64):

your process and outcome were correct, so good luck with the formatting issues

OpenStudy (anonymous):

lol thanks